# Prospects for detection of ultra high frequency gravitational waves from compact binary coalescenses with resonant cavities

### v2

@@ Abstract
This article aims at clarifying the situation about astrophysical sources that might be observed with haloscope experiments sensitive to gravitational waves in the 1-10 GHz band. The GrAHal setup is taken as a benchmark. We follow a very pedagogical path so that the full analysis can easily be used by the entire community who might not be familiar with the theoretical framework. Different relevant physical regimes are considered in details and some formulas encountered in the literature are revised. In particular, we carefully take into account the fast drift of the gravitational wave frequency and the relevant experimental timescales. We also relax the usual assumption that only the merging phase should be considered. The distances that can be probed and expected event rates are carefully evaluated, taking into account degeneracies between physical parameters. We show where experimental efforts should be focused to improve the sensitivity and we conclude that any detection in the near future is extremely unlikely.
@@ I Introduction
It is quite extraordinary that gravitational waves above the MHz might, in principle, be observed in the near future Goryachev et al. (2021); Goryachev and Tobar (2014); Berlin et al. (2022a, 2023); Domcke et al. (2022); Domcke and Garcia-Cely (2021); Herman et al. (2021a). This article aims at investigating in details the possible inspiralling binary mergers, taking the GrAHal experiment Grenet et al. (2021); T. Grenet as a benchmark. We follow a very pedagogical path so that all arguments are made clear to non-specialists. An excellent review of the possible sources can be found in Aggarwal et al. (2021) whereas an in-depth investigation of the expected signal is given in Berlin et al. (2022a). The search for light black holes was, in particular, considered in Herman et al. (2021b); Franciolini et al. (2022).
@@
Clearly, all conceivable sources in the frequency range considered here are exotic, no signal is expected from any known astrophysical source. In this article, we therefore do not focus on the a priori most favoured region of the parameter space – non is actually really favoured – but, the other way round, we try to investigate the entire range of mergers that might, in principle, be discovered. For most of the work, we remain agnostic about the formation mechanisms.
We focus on binary systems of black holes and define carefully the relevant variables. We show that a huge range of masses is a priori visible for two reasons. First, because fixing the chirp mass, which enters the evaluation of the strain, does not fix the mass of each object. Second – and more importantly – because, for a given observation frequency, varying the amount time (before coalescence) at which the system is observed changes the relevant masses. This is neglected in most studies.
After discussing in details the range of interest for the astrophysical masses, we investigate the maximum distance at which a specific source can be detected. This obviously depends not only on astrophysical parameters but also on the details of the experimental setup. We insist on introducing different important timescales and on defining relevant regimes that are overlooked in the usual literature although they play a fundamental role. Both the required strain and the maximal distance for detection are carefully studied. Our results strongly differ from the formulas often encountered in similar works, not only in numerical estimates but also in the functional dependence upon physical parameters. This is due to our “integrated” treatment of the full process – from the source to the detector –, making our output directly usable, evading the risk of evaluating things in a somewhat inconsistent way.
@@
Finally, we focus on a realistic experiment – the Grenoble Axion Haloscope (GrAHal) – and show that the orders of magnitude of distances to mergers that can be probed are very small. The resulting numbers of expected events are far from allowing any effective detection in the near future. The GrAHal project Grenet et al. (2021) is mostly designed to search for axion dark matter, but also believed to be an interesting gravitational wave detector in the ultra high frequency range. It relies on a synergy of know-hows on key aspects of such experiments: high magnetic fields and fluxes, ultra-low temperatures, and very low noise quantum detectors. Although the expected sensitivity is quite impressive, it remains insufficient for realistic astrophysical situations.
It should be underlined that quite similar questions were addressed in the pioneering work Franciolini et al. (2022). In particular, the emphasis was put on a state-of-the-art description of merger rates. It was concluded that detecting individual mergers with current proposals remains difficult. Our work is however very different as its two main ingredients were not considered in Franciolini et al. (2022). First, in the present article, we do not focus on the merging only but, the other way round, we investigate the entire parameter space opened by considering earlier stages of the coalescence. This changes the relevant masses at a given frequency by more that 10 orders of magnitude. Second, we take into account the temporal aspects more accurately by considering several regimes that are ignored in Franciolini et al. (2022). This is not a minor correction and this leads to an estimated sensitivity worsened by several orders of magnitude, therefore drastically changing the expected detection rate, even for systems observed at merging.
@@ II Observable black hole masses
More importantly, we would like to argue that the problem is not an optimization one, unless primordial black holes have a very extended and flat mass spectrum, leading to a signal spread over a wide frequency range Carr et al. (2021). We simply do not know where the sources (if any) are and the entire parameter range should therefore be investigated, even if the sensitivity is lower in some regions. There is no point focusing on specific masses or on the physical case maximising the signal as this might very well not correspond to the real World. One should keep in mind that we deal here with exotic objects whose existence and characteristics are only predicted in very specific models.
@@
In Fig.1, the chirp mass is displayed as a function of the time before merging for three of the previous configurations (the two extreme ones and the central one). For a given gravitational waves emission frequency – which means a given orbital frequency – there is a full degeneracy (ignoring for the moment other characteristics of the waves) between ${\cal M}_{c}$ and $\tau$. The chirp mass ranges from $10^{15}\text{g}\sim 10^{-18}M_{\odot}$ to $2\times 10^{-5}M_{\odot}$ for the $\nu=0.34$ GHz configuration. The upper limit becomes $2\times 10^{-6}M_{\odot}$ for the $\nu=2.67$ GHz, and $5\times 10^{-7}M_{\odot}$ for the $\nu=11.47$ GHz configuration. This covers more than 10 orders of magnitude in chirp masses.
@@
The lower bound $M_{*}\sim 10^{15}$ g simply comes from the evaporation of light black holes through the Hawking mechanism Hawking (1975). A black hole with mass $m<M_{*}$ would have fully evaporated in a time smaller than the age of the Universe. As we are confined to nearby objects, there is no chance to observe far-away lighter black holes. This is a firm limit if we assume that they were formed in the early universe. It might nevertheless make sense to also consider black holes at the Planck mass (that is 20 orders of magnitude less massive than the previously given lower bound) as quite a lot of scenarios of quantum gravity (see references in Barrau (2022)) predict that the evaporation ends with stable relics with masses $m_{rel}\sim M_{Pl}\sim 10^{-5}$ g. This would be an additional point in the plot of Fig. 1 with $\tau\sim 10^{44}$ s (way higher than the age of the universe), for the 2.67 GHz configuration. Obviously, the strain for such gravitational waves is tiny beyond words and this is a purely mathematical situation.
@@ III Distances
with $\mu_{0}$ the vacuum magnetic permeability, $\nu$ the frequency of the cavity mode $TM_{010}$, $B_{0}$ the magnetic field, $Q$ the quality factor, $V_{\text{cav}}$ the cavity volume and $\eta$ a coupling coefficient characterizing the interaction between the (dimensionless) effective current generated by a gravitational wave inside the cavity and a certain resonant mode, taken to be of order 0.1, as estimated in Berlin et al. (2022a).
@@ III.1 Conservative estimate
where $\Delta\nu$ is the spectral width, $T_{\text{sys}}$ is the system temperature (including all contributions), $k_{B}$ the Boltzmann constant and $t_{\rm eff}$ is an effective time which depends on the physical situation considered, as will be made clear in the following. It coincides with the integration time only when dealing with systems observed long before their merging.
However, it is not fully established that Eq. (7) gives the final word on the question of the SNR, and a large room for improvement can be expected, especially when considering other detection methods Lamoreaux et al. (2013). Nevertheless, we make it our baseline for calculations so as to share a common background with other studies on this topic. As it will be made clear later, it is likely that our conclusions do not depend on the detailed expression used for the SNR.
Requiring SNR $>1$ leads to [^2] In practice, it might of course be necessary to increase the SNR beyond unity. This is straightforward to implement from our analysis and this would not change the conclusion., in the quasi-circular approximation,
@@
To fix orders of magnitude, we assume that the spectral width entering this calculation is actually the bandwidth of the instrument, that is $\Delta\nu\sim\nu/Q$. This can be debated and this should be investigated in more details in the future – we discuss this point later on. This is however the correct way to go if one relies on the Dicke formula, used consistently with the signal power given by Eq. (6). Normalizing to the three sets of GrAHal fiducial values used for Fig.1, we obtain:
@@
Of course, those formulas are exactly the same ones. We chose to write 3 times the very same thing so that the orders of magnitude can be immediately obtained for fiducial configurations and unit parameters. This allows a direct and simple comparison with the estimations performed in Berlin et al. (2022a) (taking into account that we normalized to $t_{\rm eff}=1$ s and not $t_{\rm eff}=1$ min). However, the choice of the effective time $t_{\rm eff}$ is far from being straightforward and deserves a specific analysis as its expression depends upon the considered case and hides some crucial physical quantities. We will investigate this in details and will not leave it as a free parameter. This is one of the key-points of this study.
@@
The value of $t_{\Delta\nu}$ as a function of the chirp mass is shown on Fig. 3. The upper “deep” orange zone corresponds to $t_{\Delta\nu}>t_{\text{max}}\sim$ 1 year, that is longer that the full duration of the experiment whereas the upper “light” orange zone corresponds to $t_{\Delta\nu}>t_{\text{max}}\sim$ 60 s. They are extreme cases for the maximum time (respectively for the full data taking and for a single run). The lower brown zone corresponds to a time spent by the astrophysical signal in the bandwidth smaller than the inverse of the sampling frequency of the experiment, that is $t_{\Delta\nu}<t_{\text{min}}\sim Q/\nu$, as it will be explained in more details later on. Clearly, the dependence of the time spent by the signal within the instrument bandwidth upon the chirp mass is not weak and cannot be ignored.
FIG. 3: Time spent by the physical signal within the experimental bandwidth as a function of the chirp mass, for the three frequencies of interest. Upper dashed horizontal line: limit corresponding to an integration time of one year. Middle dashed horizontal line: limit corresponding to an integration time of one minute. Lower plain horizontal line: limit corresponding to $t_{\text{min}}\sim\Delta\nu^{-1}\sim 10^{-4}$ s.
@@ III.1.1 Effective time given by the signal frequency drift
The prefactor for the other setups considered here are: $2.5\times 10^{4}$ for the $0.34$ GHz configuration and $2.5\times 10^{3}$ for the $11.47$ GHz configuration (still normalizing correctly the associated experimental quantities).
Once again, we emphasize that this formula differs from what is usually encountered in the literature (not only because of prefactors but also in the exponents of the physical quantities). This is because it accounts for the fact that the time entering Eq. (16) cannot be set independently of the physical process considered. This is not a small correction but a huge effect. Importantly, the fact that lighter black holes are seen earlier in their in-spiralling history, and therefore spend more time within the detector bandwidth (as the frequency is evolving more slowly), does not counter-balance the decrease in signal intensity. As a result, the mass appears with a positive – and even larger than 1 – power in the distance evaluation.
To fix orders of magnitude, the maximum distance at which a binary system can be detected is $D_{\text{max}}\sim 0.08$ m for the lightest black holes (corresponding to $t_{\text{max}}\sim 1$ yr and black holes of mass $m=10^{-18}M_{\odot}$) and $D_{\text{max}}\sim 2.3$ AU for the heaviest ones (corresponding to $t_{\text{min}}\sim 0.1$ ms and earth-mass black holes) – one astronomical unit (AU) being of order of $10^{11}$ m. We shall comment on the numerical values (obviously tiny) in the discussion section as we keep the reasoning formal at this stage. The formula we derive can be used for arbitrary setups.
@@ III.1.2 Effective time limited by the duration of the experiment
For such small masses, the signal is very stable. This is not intrinsic to the considered systems but due to the fact that, the frequency being fixed, the inspiralling process is in this case observed long before the merging. To evaluate the SNR, one must now set $t_{\rm eff}=t_{\text{max}}$. We, however, keep $\Delta\nu=\nu/Q$. It might, in principle, be possible, in this case, to take advantage of the coherence of the signal. We will however resist the temptation of playing this game as this would spoil the easiness of comparison with other approaches and come at the price of drastically narrowing the probability that the signal falls within the experimental bandwidth. Not to mention that the physical spectral width cannot be a priori known as it depends on the unknown chirp mass.
@@
$\Leftrightarrow h$ $>$ $5.3\times 10^{-22}\times\bigg{(}\frac{2.67\ \text{GHz}}{\nu}\bigg{)}^{\frac{5}{4}}\bigg{(}\frac{0.1}{\eta}\bigg{)}\bigg{(}\frac{27\ \text{T}}{B_{0}}\bigg{)}\bigg{(}\frac{1.83\times 10^{-3}\ \text{m}^{3}}{V_{\text{cav}}}\bigg{)}^{\frac{5}{6}}\bigg{(}\frac{10^{5}}{Q}\bigg{)}^{\frac{3}{4}}\bigg{(}\frac{T_{\text{sys}}}{0.4\ \text{K}}\bigg{)}^{\frac{1}{2}}\bigg{(}\frac{60\ \text{s}}{t_{\text{max}}}\bigg{)}^{\frac{1}{4}}~{},$ (17)
@@
The prefactors for the other setups are: $0.25$ for the $0.34$ GHz configuration and also $0.25$ for the $11.47$ GHz configuration. For a haloscope experiment to observe coalescences of so light black holes, the merger would have to take place inside the resonant cavity. Beside the fact that the reasoning applied to compute the signal would clearly break down, this makes no sense. This will be shown in more details in the following sections. Interestingly, from a purely formal point of vue, the two extreme configurations lead to the same accessible distance and the 2.67 GHz is the one corresponding to the largest probed volume.
@@
Obviously, this make the situation even worst than in the previous case, as it is not even possible to take advantage of the full duration of the signal. If one sets $t_{\text{max}}\sim 1$ yr or even the time corresponding to black holes with $m=M_{*}$, the results will be catastrophic in the entire range covered by this case. (In practice 1 yr is probably not realistic as, for a very long duration, an additional pink noise should be taken into account.) If, however, the maximum integration time is, for experimental reasons, of the order of one minute, this new formula can be used to estimate accurate numbers that might, a priori not be fully out of reach.
III.1.3 Effective time limited by the sampling rate
This case corresponds to $t_{\Delta\nu}<t_{\text{min}}$. It means that the time spent by the signal within the experimental bandwidth is so small that it becomes actually smaller than the inverse sampling frequency. This also decreases the naive SNR. It is however an important situation as it corresponds to heavy masses and most of the parameter space. In this case, $t_{\rm eff}=t_{\Delta\nu}^{2}/t_{\text{min}}$, where $t_{\Delta\nu}$ is still the actual time spent by the signal in the bandwidth of the detector, given as usual by Eq. (11), and $t_{\text{min}}\sim 1/\nu_{\text{samp}}$ with $\nu_{\text{samp}}$ the sampling rate chosen at its limit, that it $\nu_{\text{samp}}\sim\Delta\nu\sim\nu/Q$. A higher sampling frequency would not improve the SNR as the cavity acts as a bandpass filter with spectral width $\Delta\nu$. We therefore use $t_{min}\sim Q/\nu$.
Requiring $SNR>1$ leads to a minimum strain for detection:

### v3

@@ Abstract
This article aims at clarifying the situation about astrophysical sources that might be observed with haloscope experiments sensitive to gravitational waves in the 1-10 GHz band. The GrAHal setup is taken as a benchmark. We follow a very pedagogical path so that the full analysis can easily be used by the entire community who might not be familiar with the theoretical framework. Different relevant physical regimes are considered in details and some approximations encountered in the literature are questioned. In particular, we carefully take into account the fast drift of the gravitational wave frequency and the associated experimental limitations of different kinds. This has strong impact on sensitivity estimates. We also relax the usual assumption that only the merging phase should be considered. The distances that can be probed are carefully evaluated, taking into account degeneracies between physical parameters. We conclude that any detection in the near future is probably impossible.
@@ I Introduction
It is quite extraordinary that gravitational waves above the MHz might, in principle, be observed thanks to resonant cavities Goryachev et al. (2021); Goryachev and Tobar (2014); Berlin et al. (2022a, 2023); Domcke et al. (2022); Domcke and Garcia-Cely (2021); Herman et al. (2021a). This article aims at investigating in details the possible inspiralling binary mergers, taking the GrAHal experiment Grenet et al. (2021); T. Grenet as a benchmark. We follow a very pedagogical path so that all arguments are made clear to non-specialists. An excellent review of the possible sources can be found in Aggarwal et al. (2021) whereas an in-depth investigation of the expected signal is given in Berlin et al. (2022a). The search for light black holes was, in particular, considered in Herman et al. (2021b); Franciolini et al. (2022).
@@
In this article, we consider binary systems of (light) black holes and define carefully the relevant variables. We do not focus on the a priori most favored region of the parameter space – non is actually really favored – but, the other way round, we try to investigate the entire range of mergers that might, in principle, be discovered. For most of the work, we remain agnostic about the formation mechanisms. We show that a huge range of masses is a priori visible for two reasons. First, because fixing the chirp mass, which enters the evaluation of the strain, does not fix the mass of each object. Second – and more importantly – because, for a given observation frequency, varying the amount time before coalescence at which the system is observed changes the relevant masses. Although neglected in most studies, this point is important because the actual mass of an hypothetical (light) binary system existing close to the Earth is not known. It could very well be different from the value corresponding to the best experimental sensitivity. Focusing on the latter case only would be relevant if the mass spectrum was known, wide, and flat – thus ensuring that the “best” situation inevitably occurs. This is not true for the kind of studies performed here: we do not know which black hole masses (if any) do exist in our cosmological neighborhood, hence the importance of considering the most general cases. As in particle phenomenology, one cannot consider the sensitivity of a detector to new particles only at the mass where the signal-to-noise ratio is the best but also at other values that might be the ones realized by Nature, even though they lead to a weaker signal.
After discussing in details the range of masses of interest, we investigate the maximum distance at which a specific source can be detected. This obviously depends not only on the astrophysical parameters but also on the details of the experimental setup. As in Franciolini et al. (2022) and Domcke et al. (2024), we insist on introducing different important timescales and on defining relevant regimes that are sometimes overlooked in the usual literature although they play a fundamental role. Both the required strain and the maximal distance for detection are carefully studied. Our results differ from the formulas sometimes encountered in similar works, not only in numerical estimates but also in the functional dependence upon physical parameters. This is due to our “integrated” treatment of the full process – from the source to the detector –, making our output directly usable, evading the risk of evaluating things in a somewhat inconsistent way.
@@
It should be underlined that quite similar questions were addressed in the pioneering work Franciolini et al. (2022). In particular, the emphasis was put on a state-of-the-art description of merger rates. It was concluded that detecting individual mergers with current proposals remains difficult. Our work is however very different as its two main ingredients were not considered in Franciolini et al. (2022). First, in the present article, we do not focus on the merging only but we investigate the entire parameter space opened by considering earlier stages of the coalescence. This changes the relevant masses at a given frequency by more that 10 orders of magnitude. This is important in itself as the actual masses are unknown but this becomes even more relevant as we show that the higher masses (corresponding to systems seen at merging) are not leading to a better sensitivity. Second, we take into account the temporal aspects more accurately by considering several regimes that are ignored in Franciolini et al. (2022). This is not a minor correction and this leads to an estimated sensitivity worsened by several orders of magnitude, therefore drastically changing the expected detection rate, even for systems observed at merging.
Numerical estimates are performed based on a realistic experiment – the Grenoble Axion Haloscope (GrAHal). The GrAHal platform Grenet et al. (2021) is mostly designed to search for axion dark matter, but was also believed to be an interesting gravitational wave detector in the ultra high frequency range. It relies on a synergy of know-hows on key aspects of such experiments: high magnetic fields and fluxes, ultra-low temperatures, and very low noise quantum detectors.
Although the expected sensitivity on the gravitational strain is quite impressive, this study shows that it remains insufficient for realistic astrophysical situations and that the orders of magnitude of distances to mergers that can be probed are very small. We conclude that binary systems of black holes are out of reach for cavities.
@@ II Observable black hole masses
More importantly, we would like to argue once more that the problem is not an optimization one, unless primordial black holes have a very extended mass spectrum, leading to a signal spread over a wide frequency range Carr et al. (2021). We simply do not know where the sources (if any) are and the entire parameter range should therefore be investigated, even if the sensitivity is lower in some regions. There is no point focusing on specific masses or on the physical case maximising the signal as this might very well not correspond to the parameters describing the real World. One should keep in mind that we deal here with exotic objects whose existence and characteristics are only predicted in very specific models. If real actual masses are such that the system frequency crosses the bandwidth of the experiment long before the merging, this situation has to be investigated as the emission at merging would anyway not be accessible to the experiment, whatever its intensity.
@@
In Fig.1, the chirp mass is displayed as a function of the time before merging for three of the previous configurations (the two extreme ones and the central one). For a given gravitational waves emission frequency – which means a given orbital frequency – there is a full degeneracy (ignoring for the moment other characteristics of the waves) between ${\cal M}_{c}$ and $\tau$. The chirp mass ranges from $10^{15}\text{g}\approx 10^{-18}M_{\odot}$ to $2\times 10^{-5}M_{\odot}$ for the $\nu=0.34$ GHz configuration. The upper limit becomes $2\times 10^{-6}M_{\odot}$ for the $\nu=2.67$ GHz, and $5\times 10^{-7}M_{\odot}$ for the $\nu=11.47$ GHz configuration. This covers more than 10 orders of magnitude in chirp masses.
@@
The lower bound $M_{*}\approx 10^{15}$ g simply comes from the evaporation of light black holes through the Hawking mechanism Hawking (1975). A black hole with mass $m<M_{*}$ would have fully evaporated in a time smaller than the age of the Universe. As we are confined to nearby objects, there is no chance to observe far-away lighter black holes. This is a firm limit if we assume that they were formed in the early universe. It might nevertheless make sense to also consider black holes at the Planck mass (that is 20 orders of magnitude less massive than the previously given lower bound) as quite a lot of scenarios of quantum gravity (see references in Barrau (2022)) predict that the evaporation ends with stable relics with masses $m_{rel}\approx M_{Pl}\approx 10^{-5}$ g. This would be an additional point in the plot of Fig. 1 with $\tau\approx 10^{44}$ s (way higher than the age of the universe), for the 2.67 GHz configuration. Obviously, the strain for such gravitational waves is tiny beyond words and this is a purely mathematical situation.
@@ III Distances
with $\mu_{0}$ the vacuum magnetic permeability, $\nu$ the frequency of the cavity mode $TM_{010}$, $B_{0}$ the magnetic field, $Q$ the cavity quality factor , $V_{\text{cav}}$ the cavity volume and $\eta$ a coupling coefficient characterizing the interaction between the (dimensionless) effective current generated by a gravitational wave inside the cavity and a certain resonant mode, taken to be of order 0.1, as estimated in Berlin et al. (2022a). In the case of very brief signals, this formula should however by modified by taking the signal quality factor instead of the cavity one. In the following the cavity quality factor is kept and this effect is accounted for through a modification of the effective time of the signal, as detailed later on.
@@ III.1 Conservative estimate
where $\Delta\nu$ is the resolution bandwidth, $T_{\text{sys}}$ is the system temperature (including all contributions), $k_{B}$ the Boltzmann constant and $t_{\rm eff}$ is an effective time which depends on the physical situation considered, as will be made clear in the following. It coincides with the integration time only when dealing with systems observed long before their merging.
Obviously, Eq. (7) is not the final word on the question of the SNR, and a room for improvement can be expected, especially when considering other detection methods Lamoreaux et al. (2013). We nevertheless make it our baseline for calculations so as to share a common background with other studies on this topic. As it will be made clear later, our conclusions do not depend on the detailed expression used for the SNR. We therefore rely on the Dicke equation and push it to its limit. This is not the only way to go but this is the working hypothesis of this article.
Requiring SNR $>1$ leads [^2] In practice, it might of course be necessary to increase the SNR beyond unity. This is straightforward to implement from our analysis and this would not change the conclusion., in the quasi-circular approximation, to
@@
To fix orders of magnitude, we first assume that the spectral width entering this calculation is actually the bandwidth of the instrument, that is $\Delta\nu\sim\nu/Q$. We then consider the optimum spectral width when the signal is more coherent than the cavity. Normalizing to the three sets of GrAHal fiducial values used for Fig.1, we obtain:
@@
Of course, those formulas are exactly the same ones. We chose to write times the very same thing so that the orders of magnitude can be immediately obtained for fiducial configurations and unit parameters. Of course, those formulas are exactly the same ones. We chose to write three times the very same thing so that the orders of magnitude can be immediately obtained for fiducial configurations and unit parameters. This allows a direct and simple comparison with the estimations performed in Berlin et al. (2022a) (taking into account that we normalized to $t_{\rm eff}=1$ s and not $t_{\rm eff}=1$ min). However, the choice of the effective time $t_{\rm eff}$ is far from being straightforward and deserves a specific analysis as its expression depends upon the considered case and hides some crucial physical quantities. We will investigate this in details and will not leave it as a free parameter. This is one of the key-points of this study.
@@
The value of $t_{\Delta\nu}$ as a function of the chirp mass is shown on Fig. 3. The upper “deep” orange zone corresponds to $t_{\Delta\nu}>t_{\text{max}}\approx$ 1 year, that is longer that the full duration of the experiment whereas the upper “light” orange zone corresponds to $t_{\Delta\nu}>t_{\text{max}}\approx$ 60 s. They are extreme cases for the maximum time (respectively for the full data taking and for a single run). The lower brown zone corresponds to a time spent by the astrophysical signal in the bandwidth smaller than the inverse of the sampling frequency of the experiment, that is $t_{\Delta\nu}<t_{\text{min}}\sim Q/\nu$, as it will be explained in more details later on. Clearly, the dependence of the time spent by the signal within the instrument bandwidth upon the chirp mass is not weak and cannot be ignored.
FIG. 3: Time spent by the physical signal within the experimental bandwidth as a function of the chirp mass, for the three frequencies of interest, with a quality factor $Q=10^{5}$. Upper dashed horizontal line: limit corresponding to an integration time of one year. Middle dashed horizontal line: limit corresponding to an integration time of one minute. Lower plain horizontal line: limit corresponding to $t_{\text{min}}\sim\Delta\nu^{-1}\approx 10^{-4}$ s.
@@
We now estimate in detail the minimum detectable strain and the maximum accessible distance for all physical cases. To keep the formalism rigorous and to make every step as clear as possible, we derive very general formulas and comment on the (obviously tiny) numerical values only in the discussion section.
@@ III.1.1 Effective time given by the signal frequency drift
The prefactor for the other setups considered here are: $2.5\times 10^{4}~{}\text{m}$ for the $0.34$ GHz configuration and $2.5\times 10^{3}~{}\text{m}$ for the $11.47$ GHz configuration (still normalizing correctly the associated experimental quantities).
Once again, we emphasize that this formula differs from what is sometimes encountered in the literature (not only because of prefactors but also in the exponents of the physical quantities). This is because it accounts for the fact that the time entering Eq. (16) cannot be set independently of the physical process considered. This is not a small correction but a huge effect. Importantly, the fact that lighter black holes are seen earlier in their in-spiralling history, and therefore spend more time within the detector bandwidth (as the frequency is evolving more slowly), does not counter-balance the decrease in signal intensity. As a result, the mass appears with a positive – and even larger than 1 – power in the distance evaluation.
@@ III.1.2 Effective time limited by the duration of the experiment
For such small masses, the signal is very stable. This is not intrinsic to the considered systems but due to the fact that, the frequency being fixed, the inspiralling process is in this case observed long before the merging. To evaluate the SNR, one must now set $t_{\rm eff}=t_{\text{max}}$. At this stage we however keep $\Delta\nu=\nu/Q$, with $Q$ the cavity quality factor.
@@
$\Leftrightarrow h$ $>$ $5.3\times 10^{-22}\times\bigg{(}\frac{2.67\ \text{GHz}}{\nu}\bigg{)}^{\frac{5}{4}}\bigg{(}\frac{0.1}{\eta}\bigg{)}\bigg{(}\frac{27\ \text{T}}{B_{0}}\bigg{)}\bigg{(}\frac{1.83\times 10^{-3}\ \text{m}^{3}}{V_{\text{cav}}}\bigg{)}^{\frac{5}{6}}\bigg{(}\frac{10^{5}}{Q}\bigg{)}^{\frac{3}{4}}\bigg{(}\frac{T_{\text{sys}}}{0.4\ \text{K}}\bigg{)}^{\frac{1}{2}}\bigg{(}\frac{60\ \text{s}}{t_{\text{max}}}\bigg{)}^{\frac{1}{4}}~{},$
@@
[…]
@@
Obviously, this make the situation even worse than in the previous case, as it is not even possible to take advantage of the full duration of the signal. If one sets $t_{\text{max}}\approx 1$ yr or even the time corresponding to black holes with $m=M_{*}$, the results will be catastrophic in the entire range covered by this case. (In practice 1 yr is probably not realistic as, for a very long duration, an additional pink noise should be taken into account.) If, however, the maximum integration time is, for experimental reasons, of the order of one minute, this new formula can be used to estimate accurate numbers that might, a priori not be fully out of reach.
III.1.3 Effective time limited by $t_{\Delta\nu}$
This is the most interesting regime as it corresponds to chirp masses $\mathcal{M}_{C}>\mathcal{M}^{\Delta\nu}_{\text{max}}$, that is to a vast portion of the parameter space which includes the highest masses accessible. To properly estimate the sensitivity, different limiting criteria must be taken into account, as the Dicke approach is pushed to its limit. Those effects are discussed below and added one by one to better illustrate how they decrease the sensitivity at high masses.
i) Physics limited case
We start by ignoring the instrumental effects and we consider an ideal (and unrealistic) case where the limit is only set by the physical drift in time of the signal frequency. We recall that even this obvious effect was sometimes neglected. This is exactly the same situation as in the previous case and the very same formulas, Eqs. (14) and (15), can be used. We simply allow them to be pushed to larger masses, namely to chirp masses ${\cal M}_{c}>\mathcal{M}_{\text{max}}^{\Delta\nu}\approx 10^{-12}M_{\odot}$. We however emphasize that the chirp mass can, anyway, not be too high, otherwise the black holes would be located inside their ISCO (or would even already have merged) before emitting gravitational waves in the GHz range, as shown in Fig. 1.
ii) Short $t_{\Delta\nu}$ effect
For chirp masses such that $\mathcal{M}_{C}>\mathcal{M}^{\Delta\nu}_{\text{max}}$ the time spent by the signal in the bandwidth of the detector $t_{\Delta\nu}$, given as usual by Eq. (11), becomes shorter than $t_{\text{min}}\approx\Delta\nu^{-1}$. The signal in the cavity lasts for $t_{\text{min}}$ (cavity decay time), hence $t_{\text{eff}}=t_{\text{min}}$. However the signal is only received during a fraction $t_{\rm eff}=t_{\Delta\nu}/t_{\text{min}}$ of the full measurement time. Compared to the case of a stationary signal, the mean power is thus reduced by the same factor. The effective time entering the Dicke equation then becomes $t_{\rm eff}=t_{\Delta\nu}^{2}/t_{\text{min}}$.
Taking this effect into account and requiring $SNR>1$ leads to a minimum strain for detection: